OpenStudy (anonymous):

A bag contains 5 chocolate, 3 peanut butter, 4 oatmeal and 4 sugar cookies. If 2 cookies are selected randomly, what is the probability that they will be the same kind?

OpenStudy (anonymous):

RESPOND QUICK

OpenStudy (anonymous):

16 cookies lets do chocolate first \[\frac{ 5 }{ 16 }\] that first one will be chocolate \[\frac{ 4 }{ 16 }\] that the next one will be chocolate or \[\frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

oops did that wrong

OpenStudy (anonymous):

\[\frac{ 4 }{ 15 }\] i meant

OpenStudy (anonymous):

\[\frac{ 3 }{ 16 }\] is the first one will be peanut butter \[\frac{ 2 }{ 15 }\] that the second one will be the same

OpenStudy (anonymous):

\[\frac{ 4 }{ 16 }\] that the first one is oatmeal/sugar \[\frac{ 1 }{ 4 }\] when simplified \[\frac{ 3 }{ 15 } or \frac{ 1 }{ 5 }\] that the second one will be as well

OpenStudy (anonymous):

I still dont get what your doing

OpenStudy (anonymous):

WHATS the answer I need this urgently

OpenStudy (anonymous):

well whats your options

OpenStudy (kirbykirby):

You can get 2 chocolate, or 2 peanut butter, or 2 oatmeal, or 2 sugar \[\Large \frac{{5\choose 2}}{{16\choose 2}}+\frac{{3\choose 2}}{{16\choose 2}}+\frac{{4\choose 2}}{{16\choose 2}}+\frac{{4\choose 2}}{{16\choose 2}}\]

OpenStudy (anonymous):

a. 6.7% b. 30% c.9% d. 93.3%

OpenStudy (kirbykirby):

This doesn't give me any of these answers. I get 25/120=5/24=20.8% Even if I try it using @KlOwNlOvE 's method which gives me the same result, it's still 5/24

OpenStudy (anonymous):

I got the answer choices off my test so I think there's a problem

OpenStudy (anonymous):

(No wonder I couldn't find the answers...)

OpenStudy (kirbykirby):

hmm.. weird.

OpenStudy (anonymous):

I know right

OpenStudy (kirbykirby):

Hopefully the answer choices are wrong o.o or maybe they incorrectly worded the question, or made a typo (maybe they mean 6 instead of 5 chocolate or smtg? :S)

OpenStudy (anonymous):

about the test I got a 90% because I missed this question 9/10

OpenStudy (kirbykirby):

=( can you check it with your teacher?

OpenStudy (anonymous):

No Im kinda bummed right now cause I cant retake it but at least its better than a 80%

OpenStudy (kirbykirby):

Well I mean like can you ask the teacher what the answer is, and how they get to it? Because there is most likely a typo in this.

OpenStudy (anonymous):

But I cant cause I'm taking an online course my parents signed me up for with NO TEACHER SUPPRORT

OpenStudy (anonymous):

Thats why im asking people on open study to help

OpenStudy (kirbykirby):

ohh i see :s

OpenStudy (anonymous):

Maybe were wrong on the question and its not a typo

OpenStudy (anonymous):

I haven't got a typo in this whole course so its rare we dint get the answer

OpenStudy (kirbykirby):

I've done quite a few of these kinds of problems recently, so it feels very fresh in my head so I'm not sure what I could have done wrong. Furthermore, this type of problem here http://math.stackexchange.com/questions/294850/3-balls-drawn-from-1-urn-probability-all-same-color-with-without-replacement is very similar to what you are asked, except replace cookies for balls, and they are interested in 3 identical balls, but you it's 2 identical cookies

OpenStudy (kirbykirby):

Although I did assume without replacement.. maybe the question wants it with replacement? :S

OpenStudy (anonymous):

I guess so Say.. We were chatting for a long time almost an hour

OpenStudy (anonymous):

I gave you a medal and I fanned you

OpenStudy (kirbykirby):

lol I didn't realize how long. I have been going throgh other questions and other windows too lol, multitasking. I just don't like leaving the person who asked the question without a full clear solution :S

OpenStudy (anonymous):

Thanks for spending your time and commitment

OpenStudy (kirbykirby):

your welcome :). Good luck with the rest of your course !

OpenStudy (anonymous):

Thank you! By the way right now I'm studying Odds in Favor and Odds Against and its pretty hard

OpenStudy (kirbykirby):

ahh . Well if you have more questions, I should be able to help :) I like probability questions hehe

OpenStudy (anonymous):

Say.. Its 7:32 right now so its time for my dinner so I have to say goodbye. Maybe we will chat again someday but my course end tomorrow so I wont be in open study that often

OpenStudy (anonymous):

Bye Bye

OpenStudy (kirbykirby):

ok, bye!